You need 150 mL of a 85% alcohol solution. On hand, you have a 70% alcohol mixture and a 95% mixture. How much of each mixture will you need to obtain the desired solution?
Answer by Alan3354(69443)

You need 150 mL of a 85% alcohol solution. On hand, you have a 70% alcohol mixture and a 95% mixture. How much of each mixture will you need to obtain the desired solution?
----------
s = amt of 75%
n = amt of 95%
------
s + n = 150
75s + 95n = 150*85